The sum of two acute angle is 110. The supplement of the smaller angle is 15 times the complement pf the larger angle. Find the measure of the acute angle.
Answer by mananth(16949)

if one angle is x the other angle is 110-x
180-x = 15(90-(110-x))
180-x= 15(-20+x)
180-x=-300+15x
300+180=16x
480=16x
x=30
The other angle is 110-30=80
30, 80 are the angles